help me please if you can.

Doubling Time Suppose a conservation campaign coupled
with higher rates causes the demand for electricity to
increase at only 2% per year, as it has recently. Find the
number of years before the utilities will need to double
generating capacity.

Answer by stanbon(75887)

Let current rate be c
EQUATION:
If current rate doubles it becomes 2c.
If the growth rate is 0.02 and it takes t years to double you get:
2c = c(1.02)^t
2 = 1.02^t
Take the log of both sides to get:
log2 = t*log1.02
t= [log2]/[log1.02]
t = 35 years
=============
Cheers,
Stan H.